OpenStudy (anonymous):

Find the inverse of the following function: f(x)= -2(x/6)+7

OpenStudy (anonymous):

Let\[f(x)=-2\left(\frac{x}{6}\right)+7=-\frac{1}{3}x+7.\]Then its inverse\[f^{-1}(x)=-3x+21.\]You can check it by composing both functions and checking that you obtain the identity:\[(f^{-1}\circ f)(x)=-3\left(-\frac{1}{3}x+7\right)+21=x,\]which holds.

OpenStudy (anonymous):

hmmmm i ended up with -3(x-7). what did i do wrong?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

\[-3(x-7)=-3x+21\]

OpenStudy (anonymous):

oh ok. thanks